# Rydberg Atomic Receiver: Next Frontier of Wireless Communications

### v1

@@ Abstract
The advancement of Rydberg Atomic REceiver (RARE) is driving a paradigm shift in electromagnetic (EM) wave measurement. RAREs utilize the electron transition phenomenon of highly-excited atoms to interact with EM waves, thereby enabling wireless signal detection. Operating at the quantum scale, such new receivers have the potential to breakthrough the sensitivity limit of classical receivers, sparking a revolution in physical-layer wireless communications. The objective of this paper is to offer insights into RARE-aided communication systems. We first provide a comprehensive introduction to the fundamental principles of RAREs. Then, a thorough comparison between RAREs and classical receivers is conducted in terms of the antenna size, sensitivity, coverage, and bandwidth. Subsequently, we overview the state-of-the-art design in RARE-aided wireless communications, exploring the latest progresses in frequency-division multiplexing, multiple-input-multiple-output, wireless sensing, and quantum many-body techniques. Finally, we highlight several wireless-communication related open problems as important research directions.
@@ I Introduction
The six-generation (6G) communication networks will leverage all available resources to support the explosive growth of global mobile services [1]. To achieve seamless coverage, 6G will expand beyond terrestrial networks in 1G–5G to space–air–ground integrated networks (SAGINs) [1]. Furthermore, the full spectrum, including sub-6GHz, millimeter-wave (mmWave), and terahertz (THz) bands, will be explored for attaining high-rate and multi-modal connections. These ambitious long-range and full-frequency communication visions require the next-generation receiver to exhibit exceptional sensitivity to electromagnetic (EM) waves across a wide spectrum.
@@
Since Hertz validated the existence of EM waves, classical receivers detect wireless signals primarily by using an antenna to transform incident EM waves into informative electric currents [2]. This mechanism, however, presents insurmountable limitations to physical-layer (PHY) wireless communications. Specifically, the front-end circuits of a receiver, such as amplifiers and mixers, introduce thermal noise into data currents. The noise power, -174 dBm/Hz, restricts the receiver’s sensitivity limit. In addition, an ideal dipole antenna typically has a half-wavelength size, which constrains its operating bandwidth to a narrow range around the center carrier frequency [1]. To detect signals across the sub-6GHz, mmWave, and THz bands, the receiver has to equip an array of bulky antennas and font-end circuits, each dedicated to a specific band. Consequently, these challenges constrain the ultimate performance of wireless communications.
@@
Recently, at the intersection of quantum sensing and wireless communication, the Rydberg Atomic REceiver (RARE) has emerged as a radical solution to overcome these limitations [3]. Rydberg atoms are highly-excited atoms with outer electrons far from the nucleus. Due to the large electric dipole moments, Rydberg atoms can strongly interact with incident EM waves, triggering electron transitions between resonate energy levels. As a result, RAREs can accurately detect wireless signals by monitoring changes on the atoms’ quantum states resulting from electron transitions [4]. Representing a paradigm shift in receiver architecture to embrace quantum sensing, RAREs are gaining increasing attentions due to their quantum advantages. For instance, a RARE is immune to thermal noise as it involves no front-end circuits. Hence, its sensitivity limit can far exceed those of classical counterparts [5]. Moreover, a single RARE is able to resonate with a large number of frequencies ranging from MHz to THz band because of the abundant energy levels in Rydberg atoms [6]. These quantum advantages make RAREs a promising technology to revolutionize PHY communications in diverse applications, such as weak signal detection, coverage enhancement, power saving, and full-frequency sensing.
@@
The objective of this paper is to offer a beginner’s guide to the next-generation wireless communications aided by RAREs. To this end, we first comprehensively introduce the fundamental features of Rydberg atoms and the key modules of a RARE. Then, the performance of RAREs is compared with classical receivers in antenna size, sensitivity, coverage, and bandwidth. Subsequently, we overview the state-of-the-art RARE techniques for wireless communications. The latest progresses in four important directions are covered: frequency-division-multiplexing (FDM), multiple-input-multiple-output (MIMO), wireless sensing, and quantum many-body techniques. Last, we highlight several wireless-communication related open problems as important future directions to conduct research on RAREs.
@@ II-A1 Definition of Rydberg atom
An atom consists of a positively charged nucleus orbited by negatively charged electrons. According to the Bohr model, electrons in these orbits have unique quantum states and quantized energy levels. Each orbit is labeled by three quantum numbers $n\ell_{j}$, (e.g., $n\ell_{j}=5S_{1/2}$), including a principal quantum number $(n=1,2,3,\cdots)$, an azimuthal quantum number $(\ell=S,P,D,\cdots)$, and a total angular momentum quantum number $(j=\frac{1}{2},\frac{3}{2},\frac{5}{2},\cdots)$. While the latter two numbers, $\ell$ and $j$, determine the angular momentum and orbital shape, the principal quantum number, $n$, affects the orbital radius and energy. For a higher value of $n$, the electron is farther from the nucleus and has higher energy (see Fig. 1). Then, Rydberg atom is defined as a highly-excited atom with one or more electrons that have a large principal quantum number, i.e., $n>20$ [4].
@@ II-A2 Electron transition for preparing Rydberg atom
Electrons can transit between energy levels by either absorbing or emitting photons from EM waves (or lasers) [5]. This electron-transition process occurs when the photon’s frequency resonates (or nearly resonates) with the transition frequency between energy levels, which is defined as the ratio between the energy-level spacing and Plank constant. The process facilitates the creation of Rydberg atoms via the excitation of electrons to high energy levels. This is typically accomplished by counter-propagating two lasers, namely a probe laser and a coupling laser, through a vapor cell filled with Alkali atoms (e.g., Rb and Cs), as shown in Fig. 1 [3]. The probe laser excites the outermost electrons from the ground state to a lowly-excited energy level, (e.g., $5S_{1/2}\rightarrow 5P_{3/2}$). The coupling beam further triggers the electron transition to a Rydberg state, (e.g. $5P_{3/2}\rightarrow 47S_{5/2}$), thereby creating Rydberg atoms. Customarily, the three involved energy levels, (e.g., $5S_{1/2}$, $5P_{3/2}$, $47S_{5/2}$), are denoted as |$\ket{1}$, |$\ket{2}$, and |$\ket{3}$ [4].
@@ II-B2 EIT based quantum measurement
To retrieve the wireless signal, it is essential to accurately readout the value of Rabi frequency. The standard measurement scheme is the EIT spectroscopy, developed based on the following two observations [9].
@@
These two observations suggest that by sweeping the probe-laser frequency around $f_{\ket{1}\rightarrow\ket{2}}$, two peaks can appear in the spectrum of the output probe laser, as depicted in Fig. 2. This allows the Rabi frequency to be measured from the peak interval, which is the principle of EIT-based quantum measurement.
@@ II-B3 Receive signal processing
After obtaining the value of Rabi frequency, the information modulated on ${\mathbf{E}}(t)$ can be recovered via customized signal processing implemented on traditional digital processors. As in Fig. 2, both amplitude-modulated (AM) and frequency-modulated (FM) symbols can be illustrated directly using the maximum likelihood method due to the dependence of Rabi frequency on $\left|{\bm{\mu}^{T}{\mathbf{E}}(t)}\right|$ and $\delta$. For the detection of phase-modulated (PM) symbols, a EIT-heterodyne scheme was developed in [7], which originates from the holographic-phase sensing technique in optics imaging. It introduces an auxiliary source to produce a known reference signal, ${\mathbf{E}}_{r}(t)$, which interferes with the unknown data signal, ${\mathbf{E}}(t)$, over the air. Their radio frequencies can be either identical or a few kHz apart. Consequently, the measured Rabi frequency changes from $|\bm{\mu}^{T}{\mathbf{E}}(t)|$ to $|\bm{\mu}^{T}{\mathbf{E}}(t)+\bm{\mu}^{T}{\mathbf{E}}_{r}(t)|$ (see the bottom-subfigure in Fig. 2). Accordingly, the information modulated on the global phase of ${\mathbf{E}}(t)$ can be extracted from the relative phase between ${\mathbf{E}}(t)$ and ${\mathbf{E}}_{r}(t)$ using spectrum analysis methods or phase retrieval algorithms [8].
@@
III Rydberg Atomic Receiver vs. Classical Receiver
@@ III-A Antenna Size
According to the Chu limit, a lossless dipole antenna in a classical receiver (CR) has a size of half-wavelength, $\frac{\lambda}{2}$. This wavelength-dependent size leads to bulky receivers at low frequencies, e.g., exceeding 10 cm when $f<1$ GHz. In contrast, the vapor cells (i.e., antennas of RAREs) have wavelength-independent sizes, e.g., typically fixed as $1$ cm in diameter [7]. Consequently, RAREs can have electrically small antennas, offering significant advantages especially for low-frequency applications.
@@ III-B Sensitivity Limit
The sensitivity of a receiver, $E_{\min}$, defined as the minimum detectable field strength per square root of averaging time (unit: $\rm V/m/\sqrt{Hz}$), is fundamentally limited by the internal noise generated by the receiver itself. For an ideal $\frac{\lambda}{2}$ dipole antenna of a CR, its internal noise stems from the thermal noise generated by front-end circuits, such as amplifiers and mixers. The corresponding sensitivity limit is expressed as $E_{\min}=\sqrt{\frac{P_{N}Z_{0}}{A_{e}}}$, where $P_{N}=-174{\rm dBm}/{\rm Hz}$ denotes the room-temperature thermal-noise power, $Z_{0}$ the free-space impedance, and $A_{e}$ the effective area. On the contrary, since a RARE contains no front-end circuits, its thermal noise is significantly weaker than a CR [5]. Its major internal noise is the quantum shot noise (QSN) resulting from the randomness when observing quantum states [2]. According to the quantum sensing theory, the QSN power is inversely proportional to the electric dipole moment, $\bm{\mu}$, the number of participating atoms, $N_{a}$, and the coherence time, $T_{r}$ of the EIT process. The theoretical sensitivity limit of a RARE, also called Standard Quantum Limit (SQL), is calculated as $E_{\min}=\frac{\hbar}{|\bm{\mu}|\sqrt{N_{a}T_{r}}}$ [2].
Figure 3: Sensitivity comparison between RAREs and CRs.
@@
Fig. 3 compares the sensitivity of RAREs with CRs. Representative values $T_{r}=225{\rm\mu s}$ and $N_{a}=5\times 10^{5}$ as in [3] are utilized. The dipole moment, $\bm{\mu}$, is selected from the electron transition $nD_{5/2}\leftrightarrow(n+1)P_{3/2}$ that resonates with each frequency. For a fair comparison, we also plot the sensitivity limit of a 1-cm dipole antenna calculated in [2] as it has a similar size with a vapor cell. Fig. 3 illustrates that the SQL can surpass the sensitivity limit of $\frac{\lambda}{2}$ dipole antennas by orders of magnitude. This is attributed to the large electric dipole moment of Rydberg atoms. Moreover, since the early exploitation of RARE in [9] (Oklahoma Univ.@2012), the RARE’s sensitivity has been improved by $10^{4}$ times. Recently developed RAREs have outperformed similarly sized dipole antennas and achieved sensitivities comparable to ideal $\frac{\lambda}{2}$ counterparts. The state-of-the-art design in [11] (Warsaw Univ.@2024) pushes the sensitivity down to $0.4{\rm\mu V/m/\sqrt{Hz}}$ at 13.9 GHz by using six-wave-mixing techniques to remove background noises. It is anticipated that in the near future, RAREs will surpass the ideal $\frac{\lambda}{2}$ dipole antenna and approach SQL, with multi-wave-mixing emerging as a promising approach for performance improvement.
III-C Long Coverage Range or Low Transmission Power
Given the extreme sensitivity promised by SQL, RAREs can theoretically detect much weaker signals than CRs, making it possible to significantly expand communication coverage. To see it more clearly, we plot Fig. 4 to illustrate the feasible communication distance, $d$, with respect to (w.r.t) the transmit power, $P_{t}$, for both RAREs and CRs in space-ground communication scenarios. The receive signal-to-noise ratio (SNR) is required to be greater than $10$ dB, which is calculated by the link budget equation: ${\rm SNR}=\frac{P_{t}GL}{P_{N}}$. The antenna gain $G$ is set as 30 dBi, and the path loss $L$ scales as $d^{-2.3}$ according to the tyical space-ground channel model. The noise power, $P_{N}$, is proportional to the square of receiver sensitivity limit, $E^{2}_{\min}$. The frequency is fixed as 6 GHz by the electron transition $70D_{5/2}\leftrightarrow 71P_{3/2}$. Fig. 4 presents the coverage-expansion capability of a SQL-approaching RARE. When $P_{t}=5\times 10^{5}$ mW and the bandwidth is 100 kHz, the coverage range can be extended from the medium-earth-orbit ($d\approx 7000$ km) up to the geostationary orbit ($d\approx 36000$ km). Moreover, it is clear from Fig. 4 that a RARE can save a large amount of transmit power for fixed coverage. If the coverage range is fixed at the medium-earth-orbit ($d=$5000 km) with a bandwidth of 1 MHz, the transmit power can be reduced from $450$ W down to $40$ W. This reveals the appealing applications of RAREs in long-range communications, like the SAGIN.
Figure 4: Comparison of feasible communication distance between RAREs and CRs. LEO, MEO, and GEO refer to the heights of low-earth-orbit, medium-earth-orbit, and geostationary satellites.
III-D Bandwidth
In current RARE-related experiments, the terminology “bandwidth” is abused into two concepts. First, the term “operating bandwidth” refers to the range of frequency that a receiver can listen to. Second, the concept of “instantaneous bandwidth” characterizes the highest wireless communication rate that a receiver can support, which follows the standard definition of “bandwidth” in the communication community.
III-D1 Operating bandwidth
The broad operating bandwidth is one advantage of RAREs. Specifically, the operating bandwidth of conventional dipole antennas is limited to several GHz due to their wavelength-dependent sizes. For example, a wideband dipole antenna centered at 10 GHz typically has an operating bandwidth of no more than 2.5 GHz. In contrast, Rydberg atoms possess abundant energy levels, allowing electrons to interact with various radio frequencies via the transition between different energy levels. For instance, the electron transitions $56D_{5/2}\leftrightarrow 57P_{3/2}$ and $56D_{5/2}\leftrightarrow 52F_{3/2}$ can resonate with the radio frequencies 12.01 GHz and 116.05 GHz, respectively. It has been calculated that a single RARE is capable of sensing a broad spectrum, ranging from MHz to THz bands [4].
III-D2 Instantaneous bandwidth

### v2

@@ Abstract
Rydberg Atomic REceiver (RARE) is driving a paradigm shift in electromagnetic (EM) wave measurement by harnessing the electron transition phenomenon of Rydberg atoms. Operating at the quantum scale, such receivers have the potential to breakthrough the performance limit of classic receivers, sparking a revolution in physical-layer wireless communications. The objective of this paper is to offer insights into RARE-empowered communications. We first provide a comprehensive introduction to the fundamental principles of RAREs. Then, a thorough comparison between RAREs and classic receivers is conducted in terms of the antenna size, sensitivity, bandwidth, and hardware cost. Subsequently, we overview the recent progresses in RARE-aided wireless communications, covering the frequency-division multiplexing, multiple-input-multiple-output, wireless sensing, and quantum many-body techniques. Moreover, the unique application of RARE in multiband sensing and communication is introduced. Finally, we conclude by providing promising research directions.
@@ I Introduction
The six-generation (6G) communication networks will leverage all available resources to support the explosive growth of global mobile services. To achieve seamless coverage, 6G will expand beyond terrestrial networks in 1G–5G to space–air–ground integrated networks. The full spectrum, including sub-6G, millimeter-wave (mmWave), and terahertz (THz) bands, will be explored for attaining high-rate and multi-modal connections. These ambitious visions require the next-generation receiver to exhibit exceptional sensitivity to electromagnetic (EM) waves across a wide spectrum.
@@
In current communication systems, receivers detect wireless signals by using an antenna to transform incident EM waves into electric currents [1]. This mechanism, however, presents insurmountable limitations to physical-layer (PHY) wireless communications. Firstly, the front-end circuits of a receiver, like amplifiers and mixers, introduce thermal noise into data currents, restricting the receiver’s sensitivity limit. In addition, an ideal dipole antenna typically has a half-wavelength size, which constrains its working bandwidth to a narrow range around the carrier frequency. To detect signals from sub-6G to THz, the receiver has to equip an array of bulky antennas and font-end circuits, each dedicated to a specific band. Consequently, these challenges constrain the ultimate performance of wireless communications [2].
@@
Recently, at the intersection of quantum sensing and wireless communication, Rydberg Atomic REceiver (RARE) has emerged as a radical solution to overcome these limitations [3]. Rydberg atoms are highly-excited atoms with outer electrons far from the nucleus. Due to the large electric dipole moments, Rydberg atoms can strongly interact with incident EM waves, triggering electron transitions between energy levels. As a result, RAREs can accurately detect wireless signals by monitoring changes on the atoms’ quantum states resulting from electron transitions [2]. Representing a paradigm shift in receiver architecture to embrace quantum sensing, RAREs are gaining increasing attentions due to their quantum advantages. For instance, a RARE is immune to thermal noise as it involves no front-end circuits. Hence, its sensitivity limit can far exceed those of classic counterparts [4]. Moreover, a single RARE is able to interact with a large number of frequencies ranging from Megahertz (MHz) to THz band because of the abundant energy levels in Rydberg atoms [5]. These quantum advantages make RAREs a promising technology to revolutionize PHY communications in diverse applications, such as weak signal detection, coverage enhancement, and full-frequency sensing.
@@
The objective of this paper is to offer a beginner’s guide to the next-generation wireless communications aided by RAREs. To this end, we first comprehensively introduce the fundamental features of Rydberg atoms and the key modules of a RARE. Then, the performance of RAREs is compared with classic receivers in antenna size, sensitivity, bandwidth, and hardware cost. Subsequently, we overview the state-of-the-art RARE techniques for wireless communications, covering frequency-division multiplexing (FDM), multiple-input-multiple-output (MIMO), wireless sensing, and quantum many-body techniques. Then, the potential of RARE in multiband sensing and communication services is revealed using experimental results. Last, we highlight promising future research directions on RARE from the PHY communication perspective.
@@ II-A1 Definition of Rydberg atom
An atom consists of a positively charged nucleus orbited by negatively charged electrons. Electrons in these orbits have unique quantum states and quantized energy levels. Each orbit is labeled by three quantum numbers $n\ell_{j}$, (e.g., $n\ell_{j}=5S_{1/2}$), including a principal quantum number $(n=1,2,3,\cdots)$, an azimuthal quantum number $(\ell=S,P,D,\cdots)$, and a total angular momentum quantum number $(j=\frac{1}{2},\frac{3}{2},\frac{5}{2},\cdots)$. While the latter two numbers, $\ell$ and $j$, determine the angular momentum and orbital shape, the principal quantum number, $n$, affects the orbital radius and energy. For a higher value of $n$, the electron is farther from the nucleus and has higher energy (see Fig. 1). Rydberg atom is defined as a highly-excited atom with one or more electrons that have a large principal quantum number, i.e., $n>20$.
@@ II-A2 Electron transition for preparing Rydberg atom
Electrons can transit between energy levels by absorbing photons from (or emitting photons to) EM waves [4]. This electron-transition process occurs when the photon’s frequency resonates with the transition frequency between energy levels, which is defined as the ratio between the energy-level spacing and Plank constant. The process facilitates the creation of Rydberg atoms via the excitation of electrons to high energy levels. This is typically accomplished by counter-propagating two lasers, namely a probe laser and a coupling laser, through a vapor cell filled with Alkali atoms (e.g., Rb and Cs), as shown in Fig. 1 [3]. The probe laser excites the outermost electrons from the ground state to a lowly-excited energy level, (e.g., $5S_{1/2}\rightarrow 5P_{3/2}$). The coupling beam further triggers the electron transition to a Rydberg state, (e.g. $5P_{3/2}\rightarrow 47S_{5/2}$), thereby creating Rydberg atoms. Customarily, the three involved energy levels, (e.g., $5S_{1/2}$, $5P_{3/2}$, $47S_{5/2}$), are denoted as |$\ket{1}$, |$\ket{2}$, and |$\ket{3}$ [2].
@@ II-B2 EIT based quantum measurement
To retrieve the wireless signal, it is essential to accurately readout the value of Rabi frequency. The standard measurement scheme is the EIT spectroscopy, developed on the following observations [9].
@@
These observations suggest that by sweeping the probe-laser frequency around $f_{\ket{1}\rightarrow\ket{2}}$, two peaks can appear in the spectrum of the output probe laser, as depicted in Fig. 2. This allows the Rabi frequency to be measured from the peak interval, which is the principle of EIT-based quantum measurement.
@@ II-B3 Receive signal processing
After obtaining the value of Rabi frequency, the information modulated on ${\mathbf{E}}(t)$ can be recovered via customized signal processing implemented on traditional digital processors. As in Fig. 2, both amplitude-modulated (AM) and frequency-modulated (FM) symbols can be illustrated directly using the maximum likelihood method due to the dependence of Rabi frequency on $\left|{\bm{\mu}^{T}{\mathbf{E}}(t)}\right|$ and $\delta$. For the detection of phase-modulated (PM) symbols, a heterodyne sensing scheme was developed in [6], which originates from the holographic-phase sensing technique in optics imaging. It introduces an auxiliary source to produce a known reference signal, ${\mathbf{E}}_{r}(t)$, which interferes with the unknown data signal, ${\mathbf{E}}(t)$, over the air. Their radio frequencies can be either identical or a few kHz apart. Consequently, the measured Rabi frequency changes from $|\bm{\mu}^{T}{\mathbf{E}}(t)|$ to $|\bm{\mu}^{T}{\mathbf{E}}(t)+\bm{\mu}^{T}{\mathbf{E}}_{r}(t)|$ (see the bottom-subfigure in Fig. 2). Accordingly, the information modulated on the global phase of ${\mathbf{E}}(t)$ can be extracted from the relative phase between ${\mathbf{E}}(t)$ and ${\mathbf{E}}_{r}(t)$ using spectrum analysis methods or phase retrieval algorithms [8].
@@
III Rydberg Atomic Receiver vs. Classic Receiver
@@ III-A Antenna Size
According to the Chu limit, a lossless dipole antenna in a classic receiver has a size of half-wavelength, $\frac{\lambda}{2}$. This wavelength-dependent size leads to bulky receivers at low frequencies, e.g., exceeding 10 cm when $f<1$ GHz. In contrast, the vapor cells (i.e., antennas of RAREs) have wavelength-independent sizes, e.g., typically fixed as $1$ cm in diameter [6]. Consequently, RAREs can have electrically small antennas, offering significant advantages especially for low-frequency applications.
@@ III-B Sensitivity Limit
The sensitivity of a receiver, $E_{\min}$, defined as the minimum detectable field strength per square root of averaging time (unit: $\rm V/m/\sqrt{Hz}$), is fundamentally limited by the internal noise generated by the receiver itself. For an ideal $\frac{\lambda}{2}$ dipole antenna of a classic receiver, its internal noise stems from the thermal noise generated by front-end circuits, such as amplifiers and mixers. The corresponding sensitivity limit is expressed as $E_{\min}=\sqrt{\frac{P_{N}Z_{0}}{A_{e}}}$, where $P_{N}=-174{\rm dBm}/{\rm Hz}$ denotes the room-temperature thermal-noise power, $Z_{0}$ the free-space impedance, and $A_{e}$ the effective area. On the contrary, since a RARE contains no front-end circuits, its thermal noise is significantly weaker than a classic receiver [4]. Its major internal noise is the quantum shot noise (QSN) resulting from the randomness when observing quantum states [1]. According to the quantum sensing theory, the QSN power is inversely proportional to the electric dipole moment, $\bm{\mu}$, the number of participating atoms, $N_{a}$, and the coherence time, $T_{r}$ of the EIT process. The theoretical sensitivity limit of a RARE, also called Standard Quantum Limit (SQL), is calculated as $E_{\min}=\frac{\hbar}{|\bm{\mu}|\sqrt{N_{a}T_{r}}}$ [1].
Figure 3: Sensitivity comparison between RAREs and classic receivers.
@@
Fig. 3 compares the sensitivity of RAREs with classic receivers. Representative values $T_{r}=225{\rm\mu s}$ and $N_{a}=5\times 10^{5}$ as in [3] are utilized. The dipole moment, $\bm{\mu}$, is selected from the electron transition $nD_{5/2}\leftrightarrow(n+1)P_{3/2}$ that resonates with each frequency. For a fair comparison, we also plot the sensitivity limit of a 1-cm dipole antenna calculated in [1] as it has a similar size with a vapor cell. Fig. 3 illustrates that the SQL can surpass the sensitivity limit of $\frac{\lambda}{2}$ dipole antennas by orders of magnitude. This is attributed to the large electric dipole moment of Rydberg atoms. Moreover, since the early exploitation of RARE in [9] (Oklahoma Univ.@2012), the RARE’s sensitivity has been improved by $10^{4}$ times. Recently developed RAREs have outperformed similarly sized dipole antennas and achieved sensitivities comparable to ideal $\frac{\lambda}{2}$ counterparts. The state-of-the-art design in [11] (Warsaw Univ.@2024) pushes the sensitivity down to $0.4{\rm\mu V/m/\sqrt{Hz}}$ at 13.9 GHz by using six-wave-mixing techniques to remove background noises. It is anticipated that in the near future, RAREs will surpass the ideal $\frac{\lambda}{2}$ dipole antenna and approach SQL, with multi-wave-mixing emerging as a promising approach for performance improvement.
III-C Bandwidth
In existing RARE studies, the term ”bandwidth” is often used to represent two distinct metrics. Firstly, it is utilized to define the range of frequencies that can trigger electron transitions of Rydberg atoms. This concept is commonly referred to as the “detectable frequency range” of a RARE in the literature. Secondly, it is also associated with the “instantaneous bandwidth”, which denotes the maximum data capacity that the receiver device can measure at any given moment. This term aligns with the standard definition of ”bandwidth” in the field of communication engineering. Typically, a RARE is characterized by a broad detectable frequency range but a narrow instantaneous bandwidth.
III-C1 Detectable frequency range
The detectable frequency range of a RARE can be significantly broader than that of a classic receiver. This is attributed to two key factors. Firstly, Rydberg atoms possess abundant energy levels. This property allows electrons to interact with various frequency bands, ranging from MHz to THz bands, via the transition between different energy levels [2]. For instance, transitions $56D_{5/2}\leftrightarrow 57P_{3/2}$ and $56D_{5/2}\leftrightarrow 52F_{3/2}$ can resonate with the radio frequencies 12.01 GHz and 116.05 GHz, respectively. Furthermore, some quantum techniques, like the Zeeman effect and off-resonate ac Stark shift, can broaden the linewidth of electron transition. In this context, the detectable frequency range of a RARE can be further expanded to achieve continuous-frequency detection exceeding 1 GHz.
III-C2 Instantaneous bandwidth